OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 12 interactions. If there are 4 children, there can be 39 interactions. Which recursive equation represents the pattern? an = an – 1 + 3(n – 1) an = an – 1 + 3(n – 1) an = an – 1 + (3n – 1) an = an – 1 + (n – 1)3

OpenStudy (anonymous):

Gotta think on this one

OpenStudy (anonymous):

a. an = a(n - 1) + 2(n - 1) a1 = 0 + 2*0 a2 = 0 + 2*1 = 2 fails at a2 b. an = a(n - 1) + (n - 1)^2 a1 = 0 + 0^2 = 0 a2 = 0 + 1^2 = 1 a3 = 1 + 2^2 = 5 a4 = 3 + 3^2 = 12 fails at a4

OpenStudy (anonymous):

So...it's: d. an = a(n - 1) + (2n - 1)

OpenStudy (anonymous):

Yes, she's right. That is the answer!

OpenStudy (dawnr):

can someone help?? http://openstudy.com/updates/5404b92de4b0f2ed1e146373 ://

OpenStudy (anonymous):

well no, but that's the formula method, I used a different amount of kids, but you should be able to solve now.

OpenStudy (anonymous):

I just needed to know the recursive pattern

OpenStudy (anonymous):

@aimee98 yes the formula is right :)

OpenStudy (anonymous):

Yup. Medal? :O

OpenStudy (anonymous):

listen @Pollylovesschool, thanks for wanting to help but you don't know what she is talking about.

OpenStudy (anonymous):

Thanks, any other questions? ;]

OpenStudy (anonymous):

@TheJonathanP4 the formula is right...she didn't give the answer...lol!

OpenStudy (anonymous):

The yearly attendance at a local movie theater is 68,000 and grows continuously at a rate of 3.5% each year. What is the approximate attendance at the movie theater in eight years? 69,603 89,543 89,973 87,040

OpenStudy (anonymous):

Okay, this is pretty simple. What's the first step here, Johnathan?

OpenStudy (anonymous):

Hey, do you mind putting a question like that instead....it keeps mentioning me to come here?

OpenStudy (anonymous):

@TheJonathanP4 do you know the first step?